A light ray is incident on a prism in minimum deviation position and suffers a deviation of 34 degree. If the shaded half of the prism is removed off, then the same ray will suffer a deviation of?

Grade:12th pass

1 Answers

Vikas TU
3 years ago
By formula δ=(n−1)A⇒34=(n−1)A 
and in the second position δ′=(n−1)A/2
∴34/δ′=(n−1)A/(n−1)A/2 
or δ′=34/2=17∘